# Learn the Ropes, Then Trust the Wins: Self-imitation with Progressive Exploration for Agentic Reinforcement Learning

### v2

@@ 3.1 Problem Definition
Given a textual task description $x\sim p(X)$ where $p(X)$ represents the distribution over task descriptions, a LLM agent parameterized by $\theta$ interacts with the environment $E$ via multiple turns until it completes the task or exceeds the maximum number of allowed turns. Such processes can be described in the following Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$:
  • State Space $\mathcal{S}$: At time $t$, the state $\mathbf{s}_{t}$ observed by the agent include the observations from the environment, previous interaction history, and the task description.
  • Action Space $\mathcal{A}$: In the context of LLM agent, the textual action $\mathbf{a}_{t}\in\mathcal{V}^{n}$ at time $t$ can either be discrete operations (e.g., task-specific tool calling) or continuous sequence (e.g., reasoning traces up to the maximum of $n$ tokens).
  • Transition Probability Function $\mathcal{P}$: The transition from the $\mathbf{s}_{t}$ to $\mathbf{s}_{t+1}$ given the action $\mathbf{a}_{t}$ is generally environment-dependent. It can be either deterministic (e.g., code interpretation) or stochastic (e.g., gaming and web environment).
  • Reward Function $\mathcal{R}$: The reward $R_{t}\in\mathbb{R}$ of the current state $\mathbf{s}_{t}$ and the action $\mathbf{a}_{t}$ can be dense (e.g., process reward), sparse, or even only at the end of the trajectory (e.g., outcome reward).
A full episode can be denoted as $\tau=\{(\mathbf{s}_{1},\mathbf{a}_{1},R_{1}),(\mathbf{s}_{2},\mathbf{a}_{2},R_{2}),...,(\mathbf{s}_{T},\mathbf{a}_{T},R_{T})\}$ with multi-turn ($T$ times) interaction. We aim to optimize the agentic LLM as a policy model $\pi_{\theta}$ by encouraging those trajectories with higher rewards. Due to the large vocabulary size $V$, the distribution modeling of the output by policy $\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t},x)$ is complex and multiple alternative actions are potential. In practice, most of the feedback from the environment may not indicate success or failure, making process supervision quite challenging. In this case, we shape agentic LLMs via RL solely with the outcome reward.
3.2 Design of Action Space for Agentic LLMs
In this study, we investigate the optimization of agentic LLMs that are capable of taking actions or using external tools to solve tasks. Building on the experimental settings proposed in recent work on agentic RL [65, 41, 34, 30], we define three distinct types of action spaces tailored to different task scenarios and use them to evaluate the effectiveness of SPEAR:
  • TextWorld Embodied Tool: The embodied actions follows ALFWorld [5] where a language-driven agent interacts with the TextWorld [66]. It allows the agent to take one of the following high-level actions: goto {recep}, take {obj} from {recep}, put {obj} in/on {recep}, open {recep}, close {recep}, toggle {obj}{recep}, clean {obj} with {recep}, heat {obj} with {recep}, and cool {obj} with {recep}, where {obj} and {recep} denote objects and receptacles, respectively.
  • Web Browsing Tool: The definition of web browsing follows WebShop [42] where only two actions are allowed: search[query] and choose[button] where query and button respectively stand for searching query and clickable elements such as back to search, prev/next page, {product title}, {option}, {desc/overview}, previous, and buy.
  • Code Interpreter Tool: The code interpreter executes the code generated by the language model and return both the stdout and stderr. If the code runs correctly, the stdout contains the output. On the other hand, the compiler error messages are provided for the next-round correction. We follow [34] to deploy a SandBox [67] service that receives execution requests from the interpreter tool. In addition, we add a reminder in the stdout for empty output when the LLM forgets to print computation results: Empty stdout! You might forget to print the answer. For non-empty stderr, we also add an instruction as hint: Errors occurred! Check your code.
@@ 3.3 Policy Optimization
Without loss of generality, we adopt the GRPO [39, 2] algorithm framework for optimizing $\pi_{\theta}$. It stems from the PPO [20] but goes one step further by replacing the model-based advantage estimation A{A} (i.e., Generalized Advantage Estimation [68]) with the intra-group relative advantage $\hat{A}$ over the “averaged” baseline.
Proximal Policy Optimization (PPO)
It maximizes the objective [69, 20]:
Group Relative Policy Optimization (GRPO)
@@ 4 Training Agentic LLMs with SPEAR [Uncaptioned image]
While the self-imitation learning (SIL) technique has been proposed as a powerful tool to train the gaming and embodied agents [37, 72, 38], its extension to the LLM-driven agents faces critical challenges of entropy collapse. Figure 3 illustrates such challenges in which the overfitting of the few available successful experiences causes irreversible stagnation of agent exploration. In addition, our preliminary experiments demonstrate that the inclusion of the auxiliary tool-call reward is a double-edged sword (as shown in Figure 5), where the competition between the reward terms causes the oscillations to converge.
To address these challenges, we introduce SPEAR for progressive exploration with self-imitation. The approach begins with skill-level exploration, where the agent is guided by a tool-call intrinsic reward to broadly investigate tool usage. Successful experiences with tool calls are then used for self-imitation, enabling action-level exploration that refines specific tool-call behaviors. This curriculum design gradually regulates policy entropy, ensuring both exploration and stability in policy learning. In addition, we extend the standard self-imitation framework with two modifications: advantage recalibration to improve off-policy estimation, and entropy-regularization to prevent policy collapse. Further details on intrinsic reward shaping, along with the proposed strong baseline Dr.BoT, are presented later in this section. For clarity, we also provide a complete summary of the training algorithm in Appendix A.1.
@@ 4.1 Self-Imitation Learning
Prioritized Experience Replay in Self-Imitation
To cope with the sparse reward problem inherent from the long-horizon tasks, we propose to leverage the SIL [37] under our RL framework. It encourages the LLM agent to imitate its past preferred experiences, driving the policy model towards taking actions whose immediate or ultimate returns are unexpectedly good. To this end, we maintain a replay buffer that stores previous trajectories, their respective rewards and advantages $\mathcal{D}=\{(\tau_{j},R_{j},\hat{A}_{j})\},j=1,2,...,N_{\mathcal{D}}$ where $N_{\mathcal{D}}$ denotes the buffer size. Under the context of GRPO training, the observed returns are only available at the end of an episode, and therefore, the original discounted returns with the trainable value model involved are simply replaced by the rule-verified rewards. To exploit only good trajectories in the replay buffer, we propose to keep those with positive in-group advantages in the off-policy GRPO loss:
@@
where the indicator function $\mathbf{1}(\cdot)$ equals to 1 when the condition satisfied and 0 otherwise. It is noted that different from the vanilla GRPO, the past trajectories might not only come from the last policy $\pi_{\theta_{old}}$ but also the policy of few steps earlier. Therefore, we heuristically use $\{\pi_{\theta_{\text{old}}}\}$ to refer to multiple previous policy models that generate the history trajectories.
Advantage Recalibration for Off-Policy Estimation
The underlying challenge of applying vanilla SIL into LLM-based agentic RL comes from the off-policy return in advantage estimation [37]. That is to say, the observed return of a trajectory from the past policy becomes increasingly different from the current one, under the assumption that the agentic LLM keeps improving during iterative policy update [38, 72]. Under this assumption, vanilla SIL computes the advantage with a pointwise max with the per-state empirical return as a baseline, which can be seen as a proxy for the upper-envelope projection of the value function onto empirical returns. GRPO removes the learned value baseline by estimating the state-dependent baseline performance through its reliance on intra-group reward averaging, but this still depends on the target policy and requires extra computation resources for sampling. Therefore, we propose to recalibrate the off-policy advantage by dynamically adjusting the baseline performance of the agent so that the relative gains of past good experiences over baselines become more associated with the current policy improvement. Specifically, we maintain a First-In-First-Out (FIFO) buffer of intra-group baselines for the latest $N_{\mathcal{D}_{R}}$ trajectories $\mathcal{D}_{R}=\{\bar{R}_{j}\}_{j=1}^{\mathcal{D}_{R}}$ where $N_{\mathcal{D}_{R}}$ denotes the size of the baseline buffer. As training progresses, due to the high variance nature of agentic RL, we utilize the 50-th percentile $P_{50}(\mathcal{D}_{R})$ as a conservative but robust estimation of the policy baseline with either upward or downward trends. To bypass the inaccurate estimation of intra-group standard deviation, we follow [40] to simply remove such a term in advantage computation:
@@
It is noted that such recalibrated advantage enjoys three benefits: 1) the baseline performance correlates with the policy changes either towards better or worse; 2) the outdated experiences in the replay buffer can be further filtered out with both $\hat{A}_{j}>0$ and $\tilde{A}_{j}>0$; 3) the difficulty bias of intra-group normalization on policy update can be mitigated. In this case, the updated off-policy self-imitation objective is:
@@
Regularization with Curriculum Schedule and Covariance-based Clipping
The trade-off between exploration and exploitation has always been the core challenge in RL [73], where the policy entropy is often used as a proxy to monitor if the policy has failed to explore new possible ways in return for suboptimal reward predictability. The policy entropy quantifies the confidence inherent in the actions triggered by the LLM. Under the context of agent tasks, we measure the average sequence level token-sum entropy of the entire trajectory $\tau$ for the policy model. Given the training data batch $\mathcal{D}_{B}$, the token-sum entropy is defined as:
Through the lens of policy entropy, we observe an entropy collapsing pattern under the tool-integrated reasoning settings (i.e., ReTool [34]). As shown in Figure 3, the vanilla GRPO training without SIL first experiences a rise in entropy which indicates the pattern transition from text-based reasoning to CI-based reasoning. Due to the feedback from the CI (e.g., code execution stderr and stdout), the policy at first exhibits uncertainty about its multi-turn actions. During RL, the policy is rewarded for certain tool-using strategy, and it becomes more certain on exploiting CI for double-check and reflection. Therefore, the entropy decreases at the later stage in exchange for more predictable, steady returns. However, the vanilla application of SIL as a complementary loss causes inevitable, monotonic entropy dropping to near zero. Such entropy collapsing rules out the possibility of policy improvement as demonstrated by the stagnant performance on AIME 2025. We believe there exist two underlying reasons:
  • The mechanical imitation of the experiences at the early stage restricts the exploration of novel reasoning paradigms. The early successful trajectories in the replay buffer are often of low diversity, where only the suboptimal solutions on simple, easy problems are stored with positive advantage. The overfitting of such trajectories reduces exploration and leads to overconfidence on the existing reasoning patterns.
  • The changes of the output logits that are highly associated with the advantage gains greatly decrease the entropy in return for rewarded behavior [17, 74]. The covariance between the log probability and the advantage of action tokens is indicative for entropy control. The over-optimization on learning tokens with high covariance improves exploitation but sacrifices exploration for further improvement.
(a) Entropy (seq-mean-token-sum-norm).
(b) Accuracy on AIME 2025.
Figure 3: Effect of our self-imitation on action-level strategy exploration (Qwen2.5-32B with code interpreter [34, 46]). The vanilla experience replay technique [37] that enforces early overfitting of the few available trajectories in the buffer causes entropy collapsing and exploration shrinkage. At the beginning, the LLM agent struggles at tool-calling skills and fails to cultivate the transition of distribution towards frequent tool utilization and tool-integrated reasoning. The naive replay limits the transformation of reasoning paradigm. In contrast, our SPEAR introduces both curriculum- and covariance- based regularization into self-imitation. Its curriculum schedule with an increasing emphasis on the replay data allows easy acquisition of skills (e.g., tool-calling) at first, and stimulates strategic action plans later. The covariance clipping removes over-confident tokens, whose log probabilities are highly associated with their advantage gains, out of optimization. Our self-imitation gives promises to exploring novel strategies and achieves steady growth on AIME 2025.
We resort to regularization based on both curriculum and covariance for entropy control. For the former, we apply a cosine warm-up schedule on the SIL term under the assumption that initially the transition of distribution towards exploration of diverse actions outweighs the imitation of fixed, limited solution patterns towards outcome success (see Figure 4). For the later, we remove tokens with high covariances [17] out of loss contribution, preventing aggressive changes of log probability for advantage acquisition.
@@
$\mathcal{J}_{\text{Total}}(\pi_{\theta})=\mathcal{J}_{\text{GRPO}}(\pi_{\theta})+\gamma\cdot\tilde{\mathcal{J}}_{\text{GRPO}}^{\text{SIL-R}}(\pi_{\theta}),$ (11)
where $t_{\text{iter}}$ and $T_{\text{warm-up}}$ respectively denote the training iteration step and the total warm-up steps. The lower bound and upper bound for determining the range of high-covariance tokens are respectively represented as $\omega_{\text{lb}}$ and $\omega_{\text{ub}}$. The operation $\operatorname{Uniform}(t|\cdot,N_{\text{clip}})$ refers to the uniform sampling of tokens $t$ with high covariance until a budget of $N_{\text{clip}}$ tokens. The indices of the selected tokens for loss masking are represented as $Ind$. It is noted that such masking introduces randomness which benefits the convergence of RL. The detailed settings of $\omega_{\text{lb}}$, $\omega_{\text{ub}}$, and $N_{\text{clip}}$ are subject to both the LLM and the task. We empirically set the rounded integers of the mean covariance in the range of top $20\%$ and top $0.02\%$ respectively for $\omega_{\text{lb}}$ and $\omega_{\text{ub}}$, and set $N^{i}_{\text{clip}}=\lambda N^{i}$ with $N^{i}$ being the total number of learnable tokens of $\tau_{i}$ and $\lambda$ denoting the clipping ratio.
Figure 4: Visualization of $\gamma$ with $T_{\text{warm-up}}=200$. The weight of SIL loss gradually increases from 0 to 1 in the first $T_{\text{warm-up}}$ steps.
4.2 Intrinsic Reward Shaping
During RL, we propose to guide the agentic LLM with a composite reward that not only considers the accuracy of the final outcome but also the beneficial behaviors that are promising to achieve the goal. Without losing generality, we adopt a sparse outcome accuracy reward, a dense, continuous tool-call reward, and a simple format reward.
Outcome Reward
Tool-call Reward
Format Reward
@@ 4.2 Intrinsic Reward Shaping
Figure 5: Effect of our intrinsic reward on skill-level strategy exploration (Qwen2.5-32B with code interpreter [34, 46]). The baseline does not consider tool-calling as a rewarded behavior and its number of interaction with the environment drops quickly due to the negative feedback (e.g., bad codes that ignore import of modules, mention undefined variables, pass unexpected indentation, and forget print of intermediate results). In this case, the LLM gives up coding and degrades to text-based reasoning. The vanilla tool-call reward, despite being effective in learning tool-call skills at first, causes competition with the accuracy reward later. Due to the limited context length, the excessive tool-call turns prevents submission of the final answer and thereafter the accuracy declines immediately. We propose to apply the curriculum schedule as an intrinsic reward design where its strength decays over step to allow the agent to merely focus on the accuracy upon mastering tool-call skills. It prevents reward hacking for unnecessarily long interactions.
Dominance of Outcome Reward with Curriculum Schedule
Although many previous studies [75, 76, 77, 78, 49, 79, 80, 81, 82, 83, 84] have experimented with various design of auxiliary rewards for improving the usage (e.g., execution success rate) of tools, our preliminary experiments show that the addition of tool-call reward is a double-edged sword. As shown in Figure 5, the agentic LLM trained without the tool-call reward fails to develop tool-integrated reasoning for solving mathematic problems under the DAPO settings [46, 34]. The reason behind such degradation to text-based reasoning paradigm is that the role of tool usage in reward acquisition is not consistent at the beginning. The success rate of tool call execution is around $0.75$ initially for Qwen2.5-32B-Instruct where most of the errors are caused by coding. Since the LLM has to fill in the written code as the argument of the interpreter function, there exist three common error patterns: 1) missing import of third-party modules such as numpy or scipy; 2) reference to variables that are defined in the history interaction context but remain undefined in the current code; 3) unexpected indentation error in the while or for loops; and 4) forgetting to print all the intermediate results that end up with empty stdout. Without the tool-call reward, the LLM agent has to deal with these errors that are newly introduced by tool calling but cause detriments to the final accuracy. Therefore, it quickly gives up tool-call and turns to pure text-based reasoning. On the other hand, the enforcement of tool-call reward stimulates an increasing number of interaction turns during training, leading to over-long responses (e.g., longer than 16K tokens) that fail to parse final answers. Such excessive tool-calls are unnecessary and detrimental to improving the outcome reward where an oscillation is observed at the later stage for both the number of tool-call turns and the performance on AIME 2025. To achieve a balance between reward terms, we alleviate the competition for dominance against the outcome reward by applying the scheduled decay on the tool-call reward:
$R^{i}=R^{i}_{\text{outcome}}+\mu\cdot R^{i}_{\text{tool-call}}+R^{i}_{\text{format}},$ (20)
where the $T_{\text{decay}}$ denotes the number of decaying steps. Figure 6 illustrates the total intrinsic reward where the dominance of the outcome reward is guaranteed.
Figure 6: Visualization of the composite intrinsic reward of outcome reward, tool-call reward, and format reward ($T_{\text{decay}}=200$). The tool-call reward gradually decays from 1 to 0 in the first 200 training steps.

### v3

@@ 3.1 Problem Definition
Given a task $x\sim p(X)$ where $p(X)$ represents data distribution, an LLM agent parameterized by $\theta$ interacts with the environment $E$ until it completes the task or exceeds the max number of turns $T$. It can be modeled by Markov Decision Process (MDP) where $\mathbf{s}_{t}$, $\mathbf{a}_{t}$, and $R_{t}$ respectively denote the state, action, and reward at time $t$. Given a full episode $\tau=\{(\mathbf{s}_{1},\mathbf{a}_{1},R_{1}),(\mathbf{s}_{2},\mathbf{a}_{2},R_{2}),$...\}, we aim to optimize the agent policy $\pi_{\theta}$. Following previous studies [64, 40, 35, 31], we define three distinct types of actions (see Appendix A.2).
@@ 3.2 Policy Optimization
We adopt the GRPO [45] which stems from PPO [65, 21] but replaces the model-based advantages A{A} [66] with the group-based $\hat{A}$ (Appendix A.3).
(a) Entropy (seq-mean-token-sum-norm).
(b) Accuracy on AIME 2025.
Figure 3: Effect of our self-imitation on action-level strategy exploration (Qwen2.5-32B with code interpreter). The vanilla experience replay technique [38] that enforces early overfitting of the few available trajectories in the buffer causes entropy collapsing and exploration shrinkage. At the beginning, the LLM agent struggles at tool-calling skills and fails to cultivate the transition of distribution towards frequent tool utilization and tool-integrated reasoning. The naive replay limits the transformation of reasoning paradigm. In contrast, our SPEAR introduces both curriculum- and covariance- based regularization into self-imitation. Its curriculum schedule with an increasing emphasis on the replay data allows easy acquisition of tool-use skills at first, and stimulates strategic action plans later. The covariance clipping removes over-confident tokens, whose log probabilities are highly associated with their advantage gains, out of optimization. Our self-imitation gives promises to exploring novel strategies and achieves steady growth on AIME 2025.
@@ 4 Training Agentic LLMs with SPEAR [Uncaptioned image]
4.1 Preliminary Findings
The extension of SIL to LLM-driven agents faces entropy collapse. Figure 3 illustrates that the overfitting of the few available successful experience causes irreversible stagnation of exploration. In addition, we demonstrate that the inclusion of the tool-call reward is a double-edged sword (Figure 4), where the competition between reward terms causes the oscillations to converge. To address these challenges, we introduce SPEAR for progressive exploration with self-imitation (Algorithm 1).
@@ 4.2 Self-Imitation Learning
We resort to self-imitation to unearth past successful experience for effective action-level exploration, where the agent learns novel strategies along the promising decision path instead of random walk and bifurcation. We prevent policy entropy divergence by replaying rewarded trajectories.
Prioritized Experience Replay in Self-Imitation.
A replay buffer is maintained to store previous trajectories, their rewards and advantages $\mathcal{D}=\{(\tau_{j},R_{j},\hat{A}_{j})\},j=1,2,...,N_{\mathcal{D}}$ where $N_{\mathcal{D}}$ denotes the buffer size. To exploit only good trajectories, we keep those with positive advantages:
@@
where the indicator $\mathbf{1}(\cdot)$ equals to 1 when the condition satisfied and 0 otherwise. The past trajectories not only come from the last policy $\pi_{\theta_{old}}$ but also the policies $\{\pi_{\theta_{\text{old}}}\}$ of few steps earlier.
Advantage Recalibration for Off-Policy Estimation.
We propose to recalibrate the advantage of trajectories in the buffer to address the underlying off-policy challenge. That is to say, the observed return of a trajectory from the past policy becomes increasingly different from the current one, under the assumption that the policy keeps improving during iterations [39, 67]. Under this assumption, vanilla SIL computes the advantage with a pointwise max with the per-state empirical return as a baseline, which can be seen as a proxy for the upper-envelope projection of the value function onto empirical returns. GRPO removes the learned value baseline by estimating the state-dependent baseline performance through its reliance on intra-group reward averaging, but this still depends on the target policy and requires extra computation resources for sampling. Dynamic adjustment on the baseline performance is performed to calibrate relative gains without introducing additional computing. Specifically, we maintain a First-In-First-Out (FIFO) buffer of intra-group baselines for the latest $N_{\mathcal{D}_{R}}$ trajectories $\mathcal{D}_{R}=\{\bar{R}_{j}\}_{j=1}^{\mathcal{D}_{R}}$ where $N_{\mathcal{D}_{R}}$ denotes the size of the baseline buffer. As training progresses, due to the high variance nature of agentic RL, we utilize the 50-th percentile $P_{50}(\mathcal{D}_{R})$ as a conservative but robust estimation of the policy baseline with either upward or downward trends. To bypass the inaccurate estimation of intra-group standard deviation, we follow [47] to simply remove such a term in advantage computation:
@@
Such recalibrated advantage enjoys three benefits: 1) the baseline performance correlates with the policy change; 2) the outdated experiences can be filtered out with both $\hat{A}_{j}>0$ and $\tilde{A}_{j}>0$; 3) the difficulty bias by group normalization can be mitigated. The updated off-policy SIL objective is:
@@
Progressive Experience Utilization with Curriculum Schedule.
We perform scheduling to 1) restrict mechanical imitation of probable-yet-immature experience at an early stage, and 2) prevent consistent uncertainty about the environment states and policy actions at later stage. We apply a warm-up $\gamma$ on the SIL term under the assumption that initially the transition of distribution towards diverse actions outweighs the imitation of limited solution patterns (see Equation 13 and Figure 6(a)).
@@
$\mathcal{J}_{\text{Total}}(\pi_{\theta})=\mathcal{J}_{\text{GRPO}}(\pi_{\theta})+\gamma\cdot\tilde{\mathcal{J}}_{\text{GRPO}}^{\text{SIL}}(\pi_{\theta}).$ (5)
@@
Figure 4: Effect of our intrinsic reward on skill-level strategy exploration (Qwen2.5-32B with code interpreter). The baseline does not consider tool-calling as a rewarded behavior and its number of interaction with the environment drops quickly due to the negative feedback of bad codes. In this case, the LLM gives up coding and degrades to text-based reasoning. The vanilla tool-call reward, despite being effective in learning tool-call skills at first, causes competition with the outcome reward later. Due to the limited context length, the excessive tool-call turns prevents submission of the final answer and thereafter the accuracy declines immediately. We propose the curriculum schedule as an intrinsic reward design where its strength decays over step to allow the agent to merely focus on the accuracy with wiser actions. It prevents reward hacking for unnecessarily long interactions.
4.3 Intrinsic Reward Shaping
We resort to intrinsic reward for skill-level exploration where the agent is guided by a tool-call reward to broadly investigate tool usage. Such design not only benefits tool learning but more importantly stimulates interactions that familiarize the agent with the environment for experience accumulation.
Reward Composition.
A compound reward $R^{i}$ of each trajectory $\tau_{i}$ not only considers the final outcome but also the behaviors that are promising to achieve the goal: an outcome accuracy reward $R^{i}_{\text{outcome}}$, a continuous tool-call reward $R^{i}_{\text{tool-call}}$, and a format reward $R^{i}_{\text{format}}$ (see Appendix A.7).
Progressive Reward Modulation with Curriculum Schedule.
We regulate the contribution of tool call reward to: 1) accelerate the mastering of tool usage for quick distribution shifting towards new task settings at an early stage, and 2) prevent optimization oscillation and competition at a later stage. Although previous studies [68, 69, 70, 71, 72, 73, 74, 75] experimented with various auxiliary rewards, we show that the addition of tool-call reward is a double-edged sword. The agent trained without the tool-call reward fails to develop tool-integrated reasoning (Figure 4) due to negative tool response: 1) missing import of modules; 2) reference to undefined variables; 3) unexpected indentation error; and 4) forgetting to print results. The agent quickly gives up coding to run away from errors and turns to pure textual reasoning. On the other hand, the enforcement of tool-call reward stimulates an increasing number of interaction turns, leading to over-long responses that cause oscillation to outcome accuracy. We alleviate the competition between reward terms by scheduling the tool-call reward with $\mu$ (Equation 14 and Figure 6(b)):
$R^{i}=R^{i}_{\text{outcome}}+\mu\cdot R^{i}_{\text{tool-call}}+R^{i}_{\text{format}}.$ (6)